# A Homothetic Gauge Theory

### v2

@@ 6.1 Homothetic Harmonicity of the Gauge Potential
Starting from the homothetic Maxwell equations (36)-(37), and using $\hat{F}=\hat{d}\hat{A}$, we choose the generalized Lorenz gauge:
@@ 6.2 Wave and Laplace Equations
7 Mathematics or Physics— HGT and the Spontaneous Symmetry Breaking
The principle of local scale invariance is a long-standing concept in theoretical physics. In Weyl’s original theory, fields transform multiplicatively. HGT proposes a different, affine-style transformation: $\phi\mapsto\phi_{d}+e^{-\lambda(x)}(\phi-\phi_{d})$. This introduces a fixed-point offset field, $\phi_{d}$, fundamentally altering the symmetry’s structure. This section argues that demanding invariance under this homothetic action is a form of spontaneous symmetry breaking (SSB). Meanwhile, the dynamics of $\lambda$ and the offset fields $\phi_{d}$ remain as the open questions.
First, it is important to recall two logically distinct types of scale transformations. In the classical Weyl approach [25], one considers local rescalings of the metric and fields as:
$g_{\mu\nu}\mapsto e^{2\sigma(x)}g_{\mu\nu},\qquad\phi\mapsto e^{W\sigma(x)}\phi,$ (48)
where $W$ is the Weyl weight of the field $\phi$. A non-Weyl-invariant model can be made invariant by introducing a compensating scalar field $\tau$ that transforms as $\tau\mapsto\tau-\sigma$. By dressing other fields with powers of $e^{\tau}$, one can achieve invariance under Weyl transformations. This is commonly interpreted as a spurionic or Stückelberg restoration of symmetry. Fixing a gauge (or assigning a vacuum expectation value to $\tau$) then breaks the symmetry spontaneously and realizes it nonlinearly. [^3] We ignore quantum anomaly issues here; the discussion is entirely classical.
In HGT, an analogous procedure applies, but the transformations act differently:
$g_{\mu\nu}\mapsto e^{-2\lambda(x)}g_{\mu\nu},\qquad\phi\mapsto\phi_{d}+e^{-\lambda(x)}(\phi-\phi_{d}),$ (49)
where $\lambda$ is the logarithm of the local scale factor, and $\phi_{d}$ is an offset field obeying the same equations as $\phi$ and transforming in the same representation. By defining $\delta\phi:=\phi-\phi_{d}$, the transformation simplifies to $\delta\phi\mapsto e^{-\lambda}\delta\phi$, meaning that $\delta\phi$ behaves like a conventional weight 1-1 field under local scalings. This encapsulates the essence of HGT transformations and allows the action to retain form-invariance under local homotheties.
Spontaneous symmetry breaking [11, 33] occurs when the underlying equations or action of a theory respect a symmetry, but the chosen vacuum state does not. In gauge theories, this typically manifests through the introduction of a scalar field acquiring a nonzero vacuum expectation value (VEV), thus selecting a specific vacuum and reducing the full symmetry to a residual subgroup. Importantly, SSB does not eliminate the symmetry but conceals it: physical configurations transform nontrivially under the broken generators. In classical field theory, this principle underlies both the Higgs mechanism and geometric symmetry reductions. Within the context of HGT, SSB becomes the bridge between mathematical symmetry and emergent physical structure, as is exemplified by the generalized Maxwell equations in the sequel.
In $\S$ 5.1, we derived the source-free generalized Maxwell equations directly from the homothetic Hodge–de Rham theory. This geometry-based approach establishes the mathematical foundation of the equations. Regardless of their physical interpretation, the homothetic gauge freedom and the derived equations are of interest in their own right, with potential applications to the theory of partial differential equations and to computational physics. Here we shall derive the same equations, in the presence of the current density $J\neq 0$, from a process that closely reflects spontaneous symmetry breaking.
The starting point is an action that is not manifestly scale-invariant due to explicit dependence on a metric and field configurations. By introducing the homothetic variables $(\phi,\phi_{d})$ and employing the transformation rule in Eq. (49), the action is modified in such a way that symmetry is effectively restored at the level of its form. This mirrors the logic of Stückelberg restoration: one modifies the theory so that it is invariant under a larger symmetry group, and the breaking is realized through a fixed gauge or background choice.
This viewpoint implies that HGT treats the effect of local scale transformation as a spontaneous symmetry breaking mechanism. However, unlike traditional gauge theories with Higgs-like scalars, the dynamical origins and physical interpretations of the fields $\lambda$ and $\phi_{d}$ remain unresolved. The field $\lambda(x)$ governs local rescalings but is not yet derived from a fundamental action. Likewise, the offset field $\phi_{d}$ is introduced to enable covariant transformations, but its dynamics and role in observable quantities have not been clarified. These gaps must be addressed if HGT is to evolve into a physical rather than purely formal framework.
7.1 Example: The Generalized Maxwell Equations and Scale-symmetry Restoration
For our derivations, we recall the following relations (from $\S$ 4.2):
  1. 1.
    The block differential on $k$-forms,
    $\hat{d}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}\ :=\ \begin{pmatrix}d\alpha+b\wedge\alpha-b\wedge\beta\\[2.0pt] d\beta\end{pmatrix},\qquad b:=d\lambda,$ (50)
    and the adjoint $\hat{\delta}:=-\hat{*}\hat{d}\hat{*}$ with respect to $\tilde{\eta}$, where $\hat{*}$ is the Hodge operator on the block bundle built from $_{\tilde{\eta}}$ on each component.
  2. 2.
    The block field-strength $\hat{F}:=\hat{d}\,\hat{A}$ for a one-form doublet $\hat{A}:=\big{(}\tilde{A},A_{d}\big{)}^{\mathsf{T}}$, with upper component
    $\tilde{F}:=d\tilde{A}+b\wedge(\tilde{A}-A_{d}),\qquad F_{d}:=dA_{d}.$ (51)
Start from the standard Maxwell system on $(M,\eta)$ coupled to a conserved current $J$,
$S_{\rm Max}[A;J]\ =\ -\frac{1}{2}\int F\wedge*_{\eta}F\ +\ \int A\wedge*_{\eta}J,\qquad F:=dA,\quad d*_{\eta}J=0,$ (52)
which in $d=4$ is conformally invariant in the sourceless sector but, in the presence of sources and a fixed choice of scale, need not be Weyl-invariant in the sense of (48). To restore local scale symmetry homothetically, introduce the compensator $\lambda$, the offset potential $A_{d}$, and the doublet $\hat{A}=(\tilde{A},A_{d})^{\mathsf{T}}$ with
@@ 7.1 Example: The Generalized Maxwell Equations and Scale-symmetry Restoration
$\tilde{A}\ :=\ A_{d}+e^{-\lambda}(A-A_{d}),\qquad\text{so that}\qquad\tilde{A}-A_{d}=e^{-\lambda}(A-A_{d}).$ (53)
We then define the HGT Maxwell action by
$S_{\rm HGT}[\,\hat{A};\lambda;\hat{J}\,]\ :=\ -\frac{1}{2}\int\big{\langle}\hat{F},\hat{F}\big{\rangle}_{\tilde{\eta}}\ +\ \int\big{\langle}\hat{A},\hat{J}\big{\rangle}_{\tilde{\eta}},\qquad\hat{F}:=\hat{d}\,\hat{A}$, (54)
where $\hat{J}:=(\tilde{J},J_{d})^{\mathsf{T}}$ is a source doublet (the top slot couples to the physical sector; the bottom slot is available for regularizations/defects and can be set to zero in the simplest applications). By construction, $S_{\rm HGT}$ is exactly invariant under local homotheties (49) with $b=d\lambda$, since all derivatives are the Weyl-covariant ones and the pairing uses $\tilde{\eta}$.
Varying $\hat{A}$ yields the generalized Maxwell equations
$\boxed{\quad\hat{d}\,\hat{F}\ =\ 0,\qquad\hat{\delta}\,\hat{F}\ =\ \hat{J},\quad}$ (55)
i.e. the Weyl-covariant Bianchi identity and source equation on the block bundle. In a Weyl-covariant Lorenz gauge,
$\hat{\delta}\,\hat{A}=0\quad\Longleftrightarrow\quad\tilde{\nabla}_{\mu}\tilde{A}^{\mu}\;-\;b_{\mu}\big{(}\tilde{A}^{\mu}-A_{d}^{\mu}\big{)}\ =\ 0,$ (56)
the equations reduce to a block wave equation
$\hat{\Delta}\,\hat{A}\ =\ \hat{J},\qquad\hat{\Delta}:=\hat{d}\,\hat{\delta}+\hat{\delta}\,\hat{d}$, (57)
with $\hat{\Delta}$ the HGT Laplace–de Rham operator built from $\tilde{\eta}$ and $b=d\lambda$. Writing out only the physical (upper) slot, one obtains the compact form
$\delta_{\tilde{\eta}}\!\left[d\tilde{A}+b\wedge(\tilde{A}-A_{d})\right]\ =\ \tilde{J},\qquad d\left[d\tilde{A}+b\wedge(\tilde{A}-A_{d})\right]\ =\ 0,$ (58)
where $\delta_{\tilde{\eta}}$ is the codifferential with respect to $\tilde{\eta}=e^{-2\lambda}\eta$. When $A_{d}$ is chosen as a pure gauge or fixed reference (e.g. $A_{d}\equiv 0$) these reduce to
$\delta_{\tilde{\eta}}\!\left(d\tilde{A}+b\wedge\tilde{A}\right)=\tilde{J},\qquad d\left(d\tilde{A}+b\wedge\tilde{A}\right)=0,$ (59)
which are the Maxwell equations in an integrable Weyl geometry with scale connection $b=d\lambda$. Equations (56)–(59) make manifest that the only effect of the restoration is to (i) replace ordinary by Weyl-covariant differentials and (ii) introduce a mixing with the offset field $A_{d}$ through $\tilde{A}-A_{d}$, i.e. through the homogeneously transforming difference.
@@ 8.1 Cauchy, Dirichlet and Neumann Boundary Conditions for the Laplace Equation
From here on, we provide a general formulation, which is valid equally for both the interior manifold $\overline{\Omega_{i}}$ and the exterior manifold $\overline{\Omega_{o}}$. Therefore, we avoided repeatedly using of indices ”ii” and ”oo”.
We denote by $\delta_{n}$  [^4] Not to be confused with the codifferential in the previous sections a sequence of regularizations of the Dirac delta function $\delta_{S}$ (it can be $S_{i}$, or $S_{o}$), such that $\lim_{n\to\infty}\delta_{n}=\delta_{S}$. Then, in the generalized Laplace equation (47), we choose (see Appendix B for justification):
@@ 8.2 A Non-Singular Model for a Point Electric Charge
$\nabla^{2}\phi_{n}+\frac{\Delta\delta_{n}}{1+\delta_{n}}(\phi_{n}-\phi_{d})=0,$ (62 revisited)
@@
9 Conclusions
@@
In this work, we have developed a Homothetic Gauge Theory (HGT) that systematically promotes local scale invariance to the status of a fundamental gauge principle, placing it on the same geometric footing as the interactions of the Standard Model. The central challenge of incorporating the non-linear operation of scaling into the linear framework of gauge theory was overcome by introducing a novel fiber-wise embedding into a doubled vector bundle. This construction linearizes the homothetic transformations, allowing them to be represented by a simple matrix group acting on the extended fibers.
@@
Upon this foundation, we constructed the full machinery of a consistent gauge theory. A generalized Weyl-integrable geometry emerged, in which the spacetime metric is conformally rescaled by a dynamical scalar gauge field, $\lambda(x)$. The mathematical integrity of the theory was secured by developing a homothetic Hodge–de Rham complex, complete with generalized differential operators and a corresponding Hodge decomposition theorem. This framework ensures that the resulting field equations are well-posed and their solutions are unique under appropriate boundary conditions.
The primary physical application of HGT was demonstrated in the context of classical electromagnetism. The theory yields a set of homothetically covariant Maxwell equations that are inherently sourceless. Traditional source terms are replaced by geometric penalty terms that couple the electromagnetic field to the gradient of the scale field, $\nabla\lambda$. This unique feature allows boundary conditions to be encoded directly into the dynamics, reformulating boundary-value problems as sourceless field equations over the entire domain.
It has shown that local scale covariance is preserved by the HGT field equations while preferred scales arise from the selection of gauge/background data. In particular, the dilatation one–form $b=d\lambda$ and the offset fields $\phi_{d}$ act as order parameters: fixing $\lambda$ and $\phi_{d}$ selects solution sectors that break scale symmetry without introducing a Higgs-type potential. Thus, HGT realizes spontaneous symmetry breaking (SSB) as a property of solutions (and their data), not of the equations themselves.
Within homothetic electromagnetism (Secs. 56), the SSB viewpoint explains how physical scales enter the generalized Maxwell theory in a controlled way. The penalization framework and homothetic Lorenz gauge make the mechanism explicit: the penalization profile and the chosen $\lambda$ determine the emergent length scale, while the equations remain homothetically covariant. This solution-level SSB suggests several directions: coupling the mechanism to gravity (Cartan/Weyl-integrable interpretations), analyzing anomalies and quantization of the homothetic sector, studying dynamical selection of $\lambda$ and $\phi_{d}$ in matter-coupled models, and exploring additional interface limits beyond electrostatics. We expect these lines to sharpen the relation between mathematical scale covariance and the emergence of physical scales across HGT applications.
@@
From the computational physicist’s viewpoint, perhaps the most significant physical consequence of this framework is a natural resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source via a volume penalization method that enforces a Dirichlet condition in the limit, we constructed a non-singular electrostatic potential that is finite everywhere, including at the origin. The divergence is eliminated not by an ad-hoc modification of Maxwell’s laws, but as a direct consequence of the underlying geometry. This result suggests that what we perceive as sources may be geometric interfaces where the scale of spacetime undergoes rapid change.
@@
Finally, by showing that the entire formalism can be understood as a specific gauge fixing within the broader context of conformal Cartan geometry, we have grounded HGT in a well-established mathematical tradition. The theory presented here is not merely a mathematical reformulation but a predictive framework with the potential for broad application. Future work could involve extending this formalism to non-Abelian gauge groups or exploring its implications for gravity, where the homothetic gauge field might offer new insights into dark energy or the cosmological constant problem.

### v3

@@ 6.1 Homothetic Harmonicity of the Gauge Potential
Starting from the homothetic Maxwell equations (41), and using $\hat{F}=\hat{d}\hat{A}$, we choose the generalized Lorenz gauge:
@@ 6.2 Wave and Laplace Equations
7 Mathematics or Physics— the roles of the scaling $\lambda$ and the offset fields $\alpha_{d}$
HGT introduces two new players into the game, the homothetic scaling $\lambda$ and the offset fields $\alpha_{d}$. Here, we connect them to the known physical quantities. Thus, HGT has the potential to be a physical theory, not just a mathematical formulation.
7.1 The Scaling $\lambda$ as the CFT dilaton field
A conservative, widely accepted heuristic is that in the weak–coupling limit a quantum gauge theory concentrates on its classical vacua [26]. For abelian gauge theory (on flat space, with no sources or $\theta$–term) those vacua are flat connections. In homothetic electromagnetism this appears as the $\lambda$–contraction $\tilde{A}\to\alpha_{d}$ as $\lambda\to\infty$, selecting the vacuum flat class. Here we argue that the same $\lambda$ that drives this contraction plays the role of the CFT dilaton (i.e. the compensator for local Weyl rescalings). The case rests on three standard, complementary tests.
(i) Local RG as a spurion: $\lambda$ shifts like a dilaton. Promote the renormalization scale to a field by the replacement
$\mu\ \longrightarrow\ \mu\,e^{-\lambda(x)}\!,\qquad\text{so that}\qquad g(x)=g\!\big(\mu e^{-\lambda(x)}\big).$ (51)
Under a local Weyl rescaling with parameter $\sigma(x)$,
$g_{\mu\nu}\to e^{2\sigma}g_{\mu\nu},\qquad\lambda\to\lambda+\sigma,\qquad\Rightarrow\quad\delta_{\sigma}g(x)=-\,\sigma(x)\,\beta(g)+\mathcal{O}(\partial\sigma),$ (52)
which is precisely Osborn’s local RG rule [25]: the compensator $\lambda$ shifts by $\sigma$ and reproduces the beta–function flow of couplings. [^3] We use “dilaton” in the standard compensator sense: if $\lambda$ is made dynamical, it is the Nambu–Goldstone mode for spontaneously broken Weyl symmetry; as a background it implements Weyl as a spurionic symmetry. Thus $\lambda$ transforms exactly as a CFT dilaton.
(ii) Weyl gauging: $b=d\lambda$ is the Weyl connection. In HGT the 1–form $b:=d\lambda$ enters the reduced field strength and equations of motion as
$\tilde{F}=d\tilde{A}+b\wedge(\tilde{A}-\alpha_{d}),\qquad d\tilde{F}+b\wedge(\tilde{F}-F_{d})=0,\qquad\delta\tilde{F}-\iota_{b}(\tilde{F}-F_{d})=e^{-\lambda}J.$ (53)
Under $\lambda\to\lambda+\sigma$ one has $b\to b+d\sigma$, the characteristic transformation of a Weyl connection. Equations (53) are precisely the form obtained by Weyl-gauging Maxwell’s equations in a weight–covariant calculus (with $A$ of weight 0 and $F$ of weight 0), and the factor $e^{-\lambda}$ multiplying $J$ is the minimal compensator required to match Weyl weight. This structural match is difficult to ascribe to coincidence: it is the expected footprint of a dilaton compensator.
(iii) Anomaly matching: the dilaton Wess–Zumino coupling. Classically in d=4d{=}4 pure Maxwell is Weyl invariant; in the presence of charged matter a nonzero $\beta(g)$ produces the familiar trace anomaly [1, 5]
$T^{\mu}_{\ \mu}\;=\;\frac{\beta(g)}{2g}\,F_{\mu\nu}F^{\mu\nu}\;+\;(\text{curvature terms}).$ (54)
The compensator/dilaton encodes this via a Wess–Zumino functional whose flat–space leading term reads
$S_{\mathrm{WZ}}[\lambda]\;\supset\;-\,\int d^{4}x\;\lambda(x)\,\frac{\beta(g)}{2g}\,F_{\mu\nu}F^{\mu\nu}\;+\;\cdots,$ (55)
so that a Weyl shift $\delta_{\sigma}\lambda=\sigma$ reproduces $\delta_{\sigma}W=\int\sigma\,T^{\mu}_{\ \mu}$. Identifying the HGT scaling field with the compensator therefore ties the $\lambda$–dependence of the effective action directly to the trace anomaly, as required for a CFT dilaton.
Synthesis. Test (i) shows that treating $\lambda$ as a spacetime–dependent subtraction scale reproduces the local RG; test (ii) shows that $b=d\lambda$ acts as the Weyl connection in the homothetic Maxwell system; test (iii) shows that coupling $\lambda$ restores Weyl invariance by encoding the trace anomaly. Together these give a coherent and standard identification:
[…]
Operationally, this means: (a) $\lambda$ shifts under local Weyl as in (52); (b) $b=d\lambda$ plays the role of the Weyl gauge field in the equations; (c) when quantum matter induces $\beta(g)\neq 0$, the $\lambda$–dependence of the effective action reproduces the trace anomaly via (55). If $\lambda$ is promoted to a dynamical field, its low–energy effective action is the standard dilaton action (including the Wess–Zumino term), while in the present work we use it as a background that organizes the homothetic reduction and weak–coupling limit.
About the kinetics of $\lambda$. In the present paper we treat $\lambda$ as a background Weyl compensator (spurion): it transforms as a dilaton and enters the Weyl–gauged Maxwell sector through $b:=d\lambda$, but it is not assigned its own kinetic term or potential. Consequently, no new propagating degree of freedom is introduced here and the theory does not predict a fifth force in this formulation. The physical effects emphasized below arise from the Weyl–gauged electromagnetism at scale gradients/interfaces and from the offset sector $\alpha_{d}$. A systematic study of an extension with a propagating dilaton will be presented elsewhere.
7.2 Offset Fields as the Vacuum Expectation Values (VEVs)
A conservative, widely accepted heuristic is that in the weak–coupling limit a quantum gauge theory concentrates on its classical vacua [26]. For abelian gauge theory on a Lorentzian spacetime without sources, boundaries, or $\theta$–terms, the classical vacua are precisely the flat connections, i.e. configurations with field strength $F=dA$ vanishing. Equivalently,
$\langle F\rangle\;=\;0\qquad\Longleftrightarrow\qquad\langle A\rangle\ \text{is flat (locally pure gauge)}.$ (57)
On simply connected spacetimes this means $\langle A\rangle=d\chi$, while on general $M$ the vacuum data are the holonomy classes $\langle A\rangle]\in H^{1}_{\mathrm{dR}}(M)$, measured by Wilson loops.
In homothetic electromagnetism we relate the physical potential $A$ and the homothetically reduced potential $\tilde{A}$ by the affine map
$\tilde{A}\;=\;e^{-\lambda}A+\bigl(1-e^{-\lambda}\bigr)\,\alpha_{d},$ (58)
so that $\lambda>0$ generates a contraction with unique fixed point $\alpha_{d}$ and $\lambda\to\infty$ implies $\tilde{A}\to\alpha_{d}$ pointwise. The corresponding field strengths are
$\tilde{F}\;=\;d\tilde{A}+b\wedge(\tilde{A}-\alpha_{d}),\qquad F_{d}\;=\;d\alpha_{d},\qquad b:=d\lambda.$ (59)
The homothetic Maxwell system (with the codifferential taken using the background metric) can be written compactly as
$d\tilde{F}+b\wedge(\tilde{F}-F_{d})=0,\qquad\delta\tilde{F}-\iota_{b}(\tilde{F}-F_{d})=e^{-\lambda}J,$ (60)
[…]
@@ 7.2 Offset Fields as the Vacuum Expectation Values (VEVs)
$\lambda\to\infty:\qquad\tilde{A}-\alpha_{d}=\mathcal{O}(e^{-\lambda}),\quad\tilde{F}\to F_{d},\quad F_{d}=0\ \ (\text{on }\mathbb{R}^{1,3}).$ (61)
The strictly gauge–invariant content of a vacuum for an abelian 1–form is captured by Wilson loops [33]. With the background–field split $A=\alpha_{d}+a$ and Gaussian fluctuations governed by a quadratic action $\propto\|da\|^{2}/g^{2}$, the weak–coupling limit $g\to 0$ (which HGT realizes as $\lambda\to\infty$) concentrates the path integral on $a=0$:
$\big\langle W_{C}\big\rangle\;=\;\left\langle\exp\!\Big(i\oint_{C}A\Big)\right\rangle\;\xrightarrow[g\to 0]{}\;\exp\!\Big(i\oint_{C}\alpha_{d}\Big).$ (62)
Hence $\alpha_{d}$ encodes the vacuum holonomy character; in a fixed gauge one may summarize this as $\alpha_{d}=\langle A\rangle$, but the gauge–invariant statement is about Wilson operators. The gauge rule is diagonal,
$(\tilde{A},\alpha_{d})\mapsto(\tilde{A}+d\chi,\;\alpha_{d}+d\chi),\qquad\Rightarrow\quad A\mapsto A+d\chi,$ (63)
so only the cohomology class $\alpha_{d}]\in H^{1}_{\mathrm{dR}}(M)$ is physical. On $\mathbb{R}^{1,3}$, $\alpha_{d}]=0$ and $\alpha_{d}$ can be gauged away.
proposition 1 (Weak coupling selects the offset vacuum)
Assume $J=0$ and no boundary flux. In homothetic electromagnetism, the limit $\lambda\to\infty$ (equivalently $g\to 0$) yields
$\lim_{\lambda\to\infty}\tilde{A}=\alpha_{d},\qquad d\alpha_{d}\ \text{harmonic (hence }=0\text{ on }\mathbb{R}^{1,3}\!),$
and for every closed curve $C$,
$\lim_{\lambda\to\infty}\,\langle W_{C}\rangle=\exp\!\Big(i\oint_{C}\alpha_{d}\Big).$
Thus the offset field $\alpha_{d}$ is the VEV of the vacuum holonomy; in a fixed gauge it may be identified with the VEV of $A$ itself.
Note that the identification above relies on the standard vacuum assumptions (no sources, boundaries, or $\theta$–terms, and Lorentz invariance). On manifolds with nontrivial $H^{1}(M)$, $\alpha_{d}$ labels distinct flat sectors (Aharonov–Bohm data). Turning on sources or boundary conditions fixes $\alpha_{d}]$ accordingly; the homothetic contraction remains, but the selected vacuum class may then be nontrivial.
@@ 8.1 Cauchy, Dirichlet and Neumann Boundary Conditions for the Laplace Equation
From here on, we provide a general formulation, which is valid equally for both the interior manifold $\overline{\Omega_{i}}$ and the exterior manifold $\overline{\Omega_{o}}$. Therefore, we avoided repeatedly using of indices "ii" and "oo".
We denote by $\delta_{n}$  [^4] Not to be confused with the codifferential in the previous sections a sequence of regularizations of the Dirac delta function $\delta_{S}$ (it can be $S_{i}$, or $S_{o}$), such that $\lim_{n\to\infty}\delta_{n}=\delta_{S}$. Then, in the generalized Laplace equation (50), we choose (see B for justification):
@@ 8.2 A Non-Singular Model for a Point Electric Charge
$\nabla^{2}\phi_{n}+\frac{\Delta\delta_{n}}{1+\delta_{n}}(\phi_{n}-\phi_{d})=0,$ (66 revisited)
@@
Remark 4 (Compatibility with a propagating $\lambda$)
Nothing in the above construction prevents promoting $\lambda$ to a propagating field in a future extension. Adding
$\mathcal{L}_{\lambda}=\tfrac{f^{2}}{2}(\partial\lambda)^{2}-V(\lambda)$
has not any conflict with the above non–singular point–charge setup because one can either:
  • (i)
    choose $V$ so that the profile used here is a stationary solution of the coupled equations (the “potential reconstruction” along the background).
  • or
  • (ii)
    pin the near–core behavior with a localized penalty term, leaving the exterior solution unchanged.
For typical regularizing profiles with $\lambda\sim a\ln r$ as $r\to 0$, the dilaton gradient energy density scales as $|\nabla\lambda|^{2}\sim a^{2}/r^{2}$, yielding a finite core contribution
$\int_{0}^{\epsilon}\!dr\,4\pi r^{2}\,|\nabla\lambda|^{2}<\infty.$
9 Conclusions and Outlook
@@
In this work, we’ve developed a Homothetic Gauge Theory (HGT) that promotes local scale invariance to a fundamental gauge principle, placing it on a geometric footing similar to the interactions of the Standard Model. We overcame the primary challenge—incorporating the non-linear action of scaling into a linear gauge framework—by embedding the tangent and cotangent bundles into doubled vector bundles. This novel construction linearizes the homothetic transformations, allowing them to be represented by a matrix group acting on the extended fibers.
@@
We then built the full machinery of a consistent gauge theory upon this foundation. A generalized Weyl-integrable geometry naturally emerged, where the spacetime metric is conformally rescaled by a scalar gauge field, $\lambda(x)$. To ensure the theory’s mathematical integrity, we developed a homothetic Hodge-de Rham complex, complete with generalized differential operators and a corresponding Hodge decomposition theorem. This framework guarantees that the resulting field equations are well-posed and that their solutions are unique under appropriate conditions.
A crucial step was to connect this mathematical formalism to established physics. We’ve shown that the scaling field $\lambda$ can be consistently identified with the conformal field theory (CFT) dilaton, serving as the compensator for local Weyl rescalings. Similarly, we’ve argued that the offset fields $\alpha_{d}$ play the role of vacuum expectation values (VEVs), which are dynamically selected in the weak-coupling limit ($\lambda\to\infty$). These identifications ground HGT’s new components in concrete physical concepts, linking them to the renormalization group and the vacuum structure of gauge theories.
The theory’s primary application was demonstrated with classical electromagnetism, yielding homothetically covariant Maxwell equations that are inherently sourceless. Instead of traditional sources, the equations feature geometric penalty terms that encode boundary conditions directly into the dynamics. This structure provides a natural mechanism for spontaneous symmetry breaking (SSB). The field equations themselves remain scale-covariant, but preferred physical scales emerge through the selection of a specific solution—a choice fixed by the background fields $\lambda$ and $\alpha_{d}$, which act as order parameters without needing a Higgs-type potential.
@@
Perhaps the most significant physical consequence of this framework is a compelling resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition on a shrinking interface, we constructed a non-singular electrostatic potential. The divergence is resolved not by an ad-hoc modification of Maxwell’s laws but as a direct result of this geometric regularization, which reinterprets sources as interfaces where the spacetime scale changes rapidly.
@@
Finally, by showing that our formalism can be understood as a specific gauge fixing within the broader context of conformal Cartan geometry, we’ve grounded HGT in a well-established mathematical tradition. The theory presented here is therefore not merely a mathematical reformulation but a predictive framework with the potential for broad application. Future work could involve extending this formalism to non-Abelian gauge groups or exploring its implications for gravity, where the homothetic gauge field might offer new insights into cosmology and fundamental physics.
Outlook: dynamical dilaton and observables. A minimal extension that we will develop in a companion paper promotes $\lambda$ to a propagating dilaton by adding
$\mathcal{L}_{\lambda}=\tfrac{f^{2}}{2}(\partial\lambda)^{2}-V(\lambda),\qquad\mathcal{L}_{\text{anom}}=\frac{\beta(e)}{2e}\,\frac{\lambda}{f}\,F_{\mu\nu}F^{\mu\nu},$
and (spurion) matter couplings at leading order $({1}/{f})\,\lambda\,T^{\mu}{}_{\mu}$. This yields:
  • (i)
    A Yukawa modification of gravity between nonrelativistic bodies,
    $V(r)=-\frac{Gm_{1}m_{2}}{r}\Big[1+\alpha\,e^{-m_{\lambda}r}\Big],\quad\alpha=2\Big(\frac{c_{m}M_{\rm Pl}}{f}\Big)^{\!2},\ \ m_{\lambda}^{2}=V^{\prime\prime}(\lambda_{0}).$
  • (ii)
    scalar–photon mixing via the anomaly vertex, alongside the HGT–specific scale–interface electromagnetic signatures already emphasized here.
The present paper confines itself to the non–propagating (compensator) regime and thus makes no fifth–force claim; the dynamical regime, parameterized by $(m_{\lambda},f)$ and the sectoral spurion weights (e.g. $c_{m}$) together with $g_{\phi\gamma}\!\sim\!\beta(e)/(ef)$, will be analyzed quantitatively in the follow–up work.